The position of a simple harmonic oscillator as a function of time is given by x=4.6cos(7πt/6 + π/8) where t is in second and x is in meter. Find: (i) the frequency (ii) the velocity and acceleration at t=0.
1
Expert's answer
2020-02-10T09:27:04-0500

As per the question,

The equation of the wave "x=4.6\cos(\dfrac{7πt}{6} + π/8)"

Now comparing this the general SHM equation, "X=A cos(\omega t+\phi)"


a)

"\omega=\dfrac{7\pi}{6}"

"\Rightarrow 2\pi f=\dfrac{7\pi}{6}"

"\Rightarrow f=\dfrac{7\pi}{6\times2\pi}=\dfrac{7}{12}Hz"


b)


"x=4.6\cos(\dfrac{7πt}{6} + π/8)"

"v=\dfrac{dx}{dt}=-4.6\times \dfrac{7\pi}{6}\sin(\dfrac{7\pi t}{6}+\dfrac{\pi}{8})"

at t=0

"v=-4.6\times \dfrac{7\pi}{6}\sin(\dfrac{\pi}{8})=-11.91m/sec"


"a=\dfrac{dv}{dt}=\dfrac{d^2x}{dt^2}=-4.6\times \dfrac{7\pi}{6}\times \dfrac{7\pi}{6}\cos(\dfrac{7\pi t}{6}+\dfrac{\pi}{8})"

at t=0,

"a=-4.6\times \dfrac{7\pi}{6}\times \dfrac{7\pi}{6}\sin(\dfrac{\pi}{8})=-\dfrac{225.4\pi^2}{36}\times\dfrac{1}{\sqrt{2}}"

"a=-43.73m\/sec^2"